# G4Splat: Geometry-Guided Gaussian Splatting with Generative Prior

### v1

@@ 4.2 Results
By incorporating geometry guidance into the generative prior, our method achieves more accurate renderings in unobserved regions while producing fewer artifacts in observed regions, as illustrated in Figs.˜4 and A5 and quantified in Table˜1. In contrast, other methods that leverage generative prior exhibit notable limitations. For example, Difix3D+ attains relatively good quality in observed regions but fails to handle unobserved areas. GenFusion, See3D, and GuidedVD can hallucinate content in unobserved regions; however, their completions are blurry and compromised by severe floaters, and they even degrade the reconstruction quality of observed regions. These comparisons highlight the effectiveness of our geometry-guided generative pipeline in maintaining high fidelity across both observed and unobserved regions.
@@
As shown in Figs.˜4 and A5, Difix3D+, GenFusion, See3D and GuidedVD all suffer from severe shape–appearance ambiguities: even when the rendered views appear visually plausible, the reconstructed geometry is of poor quality. In comparison, our approach yields more accurate geometry in unobserved regions and produces smoother, floater-free reconstructions in observed regions. This improvement stems from our plane-aware geometry modeling, which provides reliable depth supervision for both observed and unobserved areas, ensuring consistency across the entire scene. Quantitatively, Table˜1 shows that our method significantly outperforms all baselines on all reconstruction metrics across multiple datasets.
@@
As demonstrated in Figs.˜1, 5, A1 and A2, our method exhibits robust performance across a wide range of scenarios, including indoor and outdoor scenes, single-view cases and unposed scenes such as YouTube videos. As shown in Table˜A1, our method consistently outperforms all baselines regardless of the number of input views. Moreover, in scenes with complex lighting, where existing baselines struggle to achieve accurate reconstruction even with dense input views, our approach leverages accurate geometry guidance to effectively suppress errors caused by significant brightness variations across viewpoints, thereby producing high-quality reconstructions, as shown in Fig.˜A2. In addition to generating smoother planar regions, our method also faithfully reconstructs non-planar geometry. The Museum in Fig.˜1 and the Cat in Fig.˜5 further illustrate its ability to handle arbitrary 3D structures.
@@ 4.3 Ablation Studies
We conduct ablation experiments on Replica dataset to evaluate the contributions of the generative prior (GP), plane-aware geometry modeling (PM), and geometry-guided generative pipeline (PP). As shown in Table˜3, we make the following key observations:
@@
    Incorporating generative prior (GP) alone improves rendering quality but yields limited gains in geometry reconstruction. LPIPS, F-Score, and NC even decline, as GP alone tends to produce averaged, blurry results for the unseen areas. This indicates that directly introducing generative prior fails to perform as expected and leads to shape–appearance ambiguities.
@@ 4.4 Discussion on Plane Representation
Our method leverages a plane representation to obtain scale-accurate geometry guidance. We adopt this representation for three main reasons. First, planar structures are common in man-made environments and often occupy a large portion of the image, such as floors, walls, and tabletops. Second, the plane representation exhibits strong generalization. By fitting global 3D planes to accurate point clouds obtained in overlapping regions of input views, we can extend reliable depth estimation to non-overlapping or even unobserved planar regions. The accurate planar depth also enables linear alignment of monocular depth maps, improving depth accuracy of non-planar regions as well. Finally, the plane representation is simple, computationally efficient, and memory-friendly. As shown in Table˜3, our method achieves high-quality reconstruction with a runtime comparable to other approaches that employ generative prior. See additional discussions in Sections˜A.2 and C.7.
@@ A.4 More Qualitative Results
Additional qualitative results are presented in Fig.˜A5. These results further demonstrate that combining geometry guidance with the generative prior effectively alleviates shape–appearance ambiguities, leading to significant improvements in both geometry reconstruction and rendering quality, particularly in regions that are unobserved by the input views.
@@ C.3 Geometry-Guided Inpainting
As shown in Fig.˜A3, training Gaussian representations with multi-view inconsistent inpainting results leads to black shadows in inconsistent regions, a phenomenon also observed in GuidedVD Zhong et al. (2025). To address this issue, GuidedVD constrains the diffusion denoising process to preserve observed regions, thereby alleviating inconsistencies near these regions. However, it remains ineffective for large missing areas (see Figs.˜4 and A5) and suffers from slow training speed (see Table˜3).
@@ C.4 Choice of Generative Models
For the generative models in our experiments, we primarily use an inpainting-based diffusion model, i.e., See3D (Ma et al., 2025). Note that our model is compatible with other models, such as the camera-controlled diffusion models like ZeroNVS (Sargent et al., 2024) or Stable Virtual Camera (Zhou et al., 2025). In such cases, the input images serve as references, while our selected novel views act as conditions. The visible masks can thus keep the visible areas unchanged while leveraging color supervision from viewpoints that observe the global 3D plane most completely. This capability is demonstrated in the single-view testing, as detailed in Section˜A.1.
@@ C.6 Evaluation Metrics
Following prior work (Guédon et al., 2025), we evaluate rendering quality using PSNR, SSIM and LPIPS computed with the VGG network (Simonyan & Zisserman, 2014).
@@
Following prior work (Yu et al., 2022; Guédon et al., 2025), we evaluate the Chamfer Distance (CD) in $cm$, F-score with a threshold of $5cm$ and Normal Consistency (NC) for 3D scene reconstruction. These metrics are defined in Table˜A2.
@@ C.7 Implementation Details
As shown in Table˜3, training a single scene from Replica room_0 with 5 input views using our method takes slightly over one hour. In addition, we evaluate a variant with a downsampled number of Gaussians, which further accelerates training while maintaining competitive performance, as reported in the Ours (DB) entry of Table˜3.
@@
Appendix D Failure Cases and Limitations
In this section, we present and analyze representative failure cases. As shown in Fig.˜A4(a), our method is partially limited by the capabilities of current video diffusion models. For example, video diffusion models such as See3D (Ma et al., 2025) often struggle to ensure that the colors in the completed regions exactly match those of the original scene. This can lead to inconsistencies when training Gaussians with the completed images, resulting in rendering outputs that do not fully align with the visible surrounding areas. Nevertheless, by introducing scale-accurate geometry supervision, our method can achieve precise geometry reconstruction even from inconsistent completions, as demonstrated in Fig.˜A4(a).
Fig.˜A4(b) illustrates another limitation: our approach struggles with heavily occluded regions, such as the chairs partially blocked by a table. Because the table and the occluded regions of the chair are in close proximity, generating a reasonable novel camera view of the occluded areas is challenging. We hypothesize that incorporating object-level prior (Ni et al., 2025; Yang et al., 2025) could help reconstruct such severely occluded regions.

### v2

@@ 4.2 Results
By incorporating geometry guidance into the generative prior, our method achieves more accurate renderings in unobserved regions while producing fewer artifacts in observed regions, as illustrated in Figs.˜4, A6 and A8 and quantified in Tables˜1 and 2. In contrast, other methods that leverage generative prior exhibit notable limitations. For example, Difix3D+ attains relatively good quality in observed regions but fails to handle unobserved areas. GenFusion, See3D, and GuidedVD can hallucinate content in unobserved regions; however, their completions are blurry and compromised by severe floaters, and they even degrade the reconstruction quality of observed regions. These comparisons highlight the effectiveness of our geometry-guided generative pipeline in maintaining high fidelity across both observed and unobserved regions.
@@
As shown in Figs.˜4, A6 and A8, Difix3D+, GenFusion, See3D and GuidedVD all suffer from severe shape–appearance ambiguities: even when the rendered views appear visually plausible, the reconstructed geometry is of poor quality. In comparison, our approach yields more accurate geometry in unobserved regions and produces smoother, floater-free reconstructions in observed regions. This improvement stems from our plane-aware geometry modeling, which provides reliable depth supervision for both observed and unobserved areas, ensuring consistency across the entire scene. Quantitatively, Table˜1 shows that our method significantly outperforms all baselines on all reconstruction metrics across multiple datasets.
@@
As demonstrated in Figs.˜1, 5, A1 and A2, our method exhibits robust performance across a wide range of scenarios, including indoor and outdoor scenes, single-view cases and unposed scenes such as YouTube videos. As shown in Table˜A1, our method consistently outperforms all baselines regardless of the number of input views. Moreover, in scenes with complex lighting, where existing baselines struggle to achieve accurate reconstruction even with dense input views, our approach leverages accurate geometry guidance to effectively suppress errors caused by significant brightness variations across viewpoints, thereby producing high-quality reconstructions, as shown in Fig.˜A2. In addition to achieving superior results on indoor scenes, as shown in the quantitative and qualitative results on the Mip-NeRF 360 dataset (Tables˜2, A3, 4 and A7), our method also outperforms the baselines on outdoor, non-Manhattan, and less structured scenes. Furthermore, the Museum in Fig.˜1 and the Cat in Fig.˜5 further demonstrate its capability to handle arbitrary 3D structures.
@@ 4.3 Ablation Studies
We conduct ablation experiments on Replica dataset to evaluate the contributions of the generative prior (GP), plane-aware geometry modeling (PM), and geometry-guided generative pipeline (PP). As shown in Tables˜4 and A4, we make the following key observations:
@@
    Incorporating generative prior (GP) alone improves rendering quality but yields limited gains in geometry reconstruction. LPIPS, F-Score, and NC even decline, as GP alone tends to produce averaged, blurry results for the unseen areas. This indicates that directly introducing generative prior fails to perform as expected and leads to shape–appearance ambiguities. Moreover, as shown in Fig.˜A4, our method is compatible with different generative diffusion models, achieving strong performance across them.
@@ 4.4 Discussion on Plane Representation
Our method leverages a plane representation to obtain scale-accurate geometry guidance. We adopt this representation for three main reasons. First, planar structures are common in man-made environments and often occupy a large portion of the image, such as floors, walls, and tabletops. Second, the plane representation exhibits strong generalization. By fitting global 3D planes to accurate point clouds obtained in overlapping regions of input views, we can extend reliable depth estimation to non-overlapping or even unobserved planar regions. The accurate planar depth also enables linear alignment of monocular depth maps, improving depth accuracy of unobserved non-planar regions as well. Finally, the plane representation is simple, computationally efficient, and memory-friendly. As shown in Table˜4, our method achieves high-quality reconstruction with a runtime comparable to other approaches that employ generative prior. Our accelerated variant, Ours (DS), which downsamples the initial Gaussians, substantially reduces runtime while still outperforming all baselines.
Furthermore, our method maintains robust performance in non-planar or less structured scenes because it is a strict enhancement of the base model. This stems from our design that applies tailored supervision on regions according to their geometric characteristics: for planar regions, reconstruction is improved by leveraging accurate plane depth supervision across both observed and unobserved areas; for non-planar regions, observed regions preserve the base model’s original supervision to ensure reliability, and unobserved portions benefit from linearly aligned monocular depth as an effective fallback where direct supervision is absent. Experiments on the Mip-NeRF 360 dataset validate these advantages. Additional discussions are provided in Sections˜A.2 and C.7.
@@ A.4 More Qualitative Results
Additional qualitative results are presented in Figs.˜A7 and A8. These results further demonstrate that combining geometry guidance with the generative prior effectively alleviates shape–appearance ambiguities, leading to significant improvements in both geometry reconstruction and rendering quality, particularly in regions that are unobserved by the input views.
@@ C.3 Geometry-Guided Inpainting
As shown in Fig.˜A3, training Gaussian representations with multi-view inconsistent inpainting results leads to black shadows in inconsistent regions, a phenomenon also observed in GuidedVD Zhong et al. (2025). To address this issue, GuidedVD constrains the diffusion denoising process to preserve observed regions, thereby alleviating inconsistencies near these regions. However, it remains ineffective for large missing areas (see Figs.˜4 and A8) and suffers from slow training speed (see Table˜4).
@@ C.4 Choice of Generative Models
For the generative models in our experiments, we primarily employ the inpainting-based diffusion model, See3D (Ma et al., 2025). Notably, our framework remains highly effective when integrated with alternative generative priors, such as ViewCrafter (Yu et al., 2024b), as evidenced by the quantitative and qualitative results in Tables˜A2 and A4. Beyond these, our model is also compatible with camera-controlled diffusion models like ZeroNVS (Sargent et al., 2024) or Stable Virtual Camera (Zhou et al., 2025). In these configurations, the input images serve as references, while our selected novel views act as conditions. The visible masks can thus keep the visible areas unchanged while leveraging color supervision from viewpoints that observe the global 3D plane most completely. This capability is demonstrated in the single-view testing, as detailed in Section˜A.1.
Table A2: Quantitative results of our method integrated with different diffusion models.
Method Reconstruction Rendering
[…]
MAtCha 10.12 60.90 79.33 17.81 0.752 0.228
GenFusion 13.05 41.60 69.33 20.14 0.801 0.258
Difix3D+ 13.71 43.11 65.34 19.42 0.779 0.231
GuidedVD 27.87 17.29 61.64 22.51 0.822 0.260
Ours (w. ViewCrafter) 7.95 64.19 81.82 22.08 0.812 0.210
Ours (w. See3D) 6.61 65.14 83.98 23.90 0.836 0.199
Figure A4: Qualitative comparison from 5 input views using different diffusion models. G4Splat produces high-quality, smooth geometry reconstruction even when paired with a weaker generative prior (i.e. Ours with ViewCrafter), and it continues to effectively suppress Gaussian floaters and reduce blurriness in the rendered results.
@@ C.6 Evaluation Metrics
Following prior work (Guédon et al., 2025), we evaluate rendering quality using PSNR, SSIM and LPIPS computed with the VGG network (Simonyan and Zisserman, 2014).
@@
Following prior work (Yu et al., 2022; Guédon et al., 2025), we evaluate the Chamfer Distance (CD) in cmcm, F-score with a threshold of $5cm$ and Normal Consistency (NC) for 3D scene reconstruction. These metrics are defined in Table˜A4.
@@ C.7 Implementation Details
As shown in Table˜4, training a single scene from Replica room_0 with 5 input views using our method takes slightly over one hour. In addition, we evaluate a variant with a downsampled number of Gaussians, which further accelerates training while maintaining competitive performance, as reported in the Ours (DS) entry of Table˜4.
Appendix D Failure Cases and Limitations
In this section, we present and analyze representative failure cases. As shown in Fig.˜A5(a), our method is partially limited by the capabilities of current video diffusion models. For example, video diffusion models such as See3D (Ma et al., 2025) often struggle to ensure that the colors in the completed regions exactly match those of the original scene. This can lead to inconsistencies when training Gaussians with the completed images, resulting in rendering outputs that do not fully align with the visible surrounding areas. Nevertheless, by introducing scale-accurate geometry supervision, our method can achieve precise geometry reconstruction even from inconsistent completions, as demonstrated in Fig.˜A5(a).
Fig.˜A5(b) illustrates another limitation: our approach struggles with heavily occluded regions, such as the chairs partially blocked by a table. Because the table and the occluded regions of the chair are in close proximity, generating a reasonable novel camera view of the occluded areas is challenging. We hypothesize that incorporating object-level prior (Ni et al., 2025; Yang et al., 2025) could help reconstruct such severely occluded regions.
@@ Appendix E LLM Usage Statement
Figure A6: Qualitative comparison from 5 input views with visibility mask. In the GT images, golden regions indicate areas unobserved across the 5 input views. G4Splat outperforms baselines in both visible and unobserved regions, producing superior geometry with improved smoothness and minimal Gaussian artifacts.
Figure A7: Qualitative comparison using 9 input views on the Mip-NeRF 360 dataset. We present results for all 5 outdoor scenes in the dataset. Our method produces significantly fewer Gaussian floaters and less blurriness, while recovering higher-quality and smoother geometry. These results demonstrate that our approach also performs robustly on non-planar and less structured outdoor scenes.
Table A3: Quantitative results from 9 input views on the Mip-NeRF 360 dataset. G4Splat performs strongly not only on indoor, Manhattan-style scenes (bonsai, counter, kitchen, room), but also on outdoor, non-Manhattan, and less structured scenes (bicycle, flowers, garden, stump, treehill), demonstrating its robustness in non-planar environments. Top-3 results are highlighted as the first, second and third.
bicycle bonsai counter flowers garden kitchen room stump treehill Mean
PSNR $\uparrow$
MAtCha 15.56 15.36 15.96 13.66 17.94 18.96 17.16 18.52 14.80 16.44
GenFusion 16.50 17.93 17.96 14.37 18.43 19.76 19.80 19.82 15.79 17.82
Difix3D+ 15.23 16.66 16.68 13.41 17.54 18.68 17.17 18.08 13.03 16.28
GuidedVD 14.89 17.21 17.60 14.23 18.23 18.00 18.50 19.00 13.36 16.78
See3D 16.08 16.32 16.49 15.80 19.94 19.62 18.23 18.26 11.51 16.92
Ours 18.31 17.97 19.94 15.48 19.18 20.84 20.71 19.53 15.99 18.66
SSIM $\uparrow$
MAtCha 0.281 0.541 0.549 0.218 0.390 0.705 0.616 0.397 0.381 0.453
GenFusion 0.296 0.586 0.577 0.237 0.339 0.562 0.667 0.422 0.383 0.452
Difix3D+ 0.227 0.551 0.497 0.185 0.288 0.530 0.568 0.339 0.313 0.389
GuidedVD 0.274 0.550 0.570 0.221 0.345 0.578 0.646 0.403 0.359 0.438
See3D 0.229 0.513 0.526 0.235 0.416 0.732 0.666 0.388 0.280 0.443
Ours 0.345 0.630 0.612 0.289 0.454 0.725 0.733 0.460 0.386 0.515
LPIPS $\downarrow$
MAtCha 0.482 0.362 0.339 0.520 0.351 0.241 0.325 0.441 0.443 0.389
GenFusion 0.530 0.384 0.384 0.568 0.429 0.335 0.319 0.493 0.513 0.439
Difix3D+ 0.509 0.370 0.378 0.529 0.408 0.311 0.350 0.474 0.501 0.426
GuidedVD 0.548 0.444 0.408 0.576 0.448 0.380 0.646 0.502 0.535 0.499
See3D 0.553 0.441 0.394 0.462 0.369 0.394 0.339 0.538 0.564 0.451
Ours 0.484 0.336 0.330 0.450 0.314 0.242 0.286 0.445 0.457 0.371